# Speak, Segment, Track, Navigate: An Interactive System for Video-Guided Skull-Base Surgery

### 17 Mar 2026 v1

@@ IV-A Pose Tracking Accuracy
TABLE I: Absolute tool tip position differences in the anatomy frame.
$|\Delta x|$ ($m$) $|\Delta y|$ ($m$) $|\Delta z|$ ($m$) $\|\Delta\mathbf{p}\|_{2}$ ($m$)
Trial 1 1.53 $\pm$ 0.67 1.37 $\pm$ 2.04 1.81 $\pm$ 0.67 2.74 $\pm$ 1.74
Trial 2 0.72 $\pm$ 0.57 2.01 $\pm$ 1.77 0.49 $\pm$ 0.52 2.32 $\pm$ 1.76
Trial 3 1.21 $\pm$ 0.48 1.02 $\pm$ 1.84 2.73 $\pm$ 0.54 3.44 $\pm$ 1.43
Mean 1.15 $\pm$ 0.57 1.47 $\pm$ 1.88 1.68 $\pm$ 0.58 2.83 $\pm$ 1.64
Note: Differences are between the optical tracker and the vision-based method, expressed in the anatomy coordinate frame. Values are mean $\pm$ standard deviation across time-matched frames for each trial.
Inter-frame rotation propagation discrepancy.
Because absolute orientation can depend on convention and initialization, we compare relative inter-frame motion. Let $\mathbf{R}^{V}_{A,i}\in SO(3)$ and $\mathbf{R}^{O}_{A,i}\in SO(3)$ denote the vision-based and optical-tracker rotations expressed in $\mathcal{F}_{A}$. We compute per-step increments
$\Delta\mathbf{R}^{(\cdot)}_{i}=\left(\mathbf{R}^{(\cdot)}_{A,i-1}\right)^{\!\top}\mathbf{R}^{(\cdot)}_{A,i}$, (4)
and define the per-step propagation discrepancy as
@@
Figure 12: Trial 1 tool-tip trajectory in the anatomy frame. The $x$, $y$, and $z$ components are plotted over time (time-matched sample index) for the vision-based method and the optical tracker.
Figure 13: Trial 2 tool-tip trajectory in the anatomy frame.
Figure 14: Trial 3 tool-tip trajectory in the anatomy frame.
TABLE II: Inter-frame relative rotation propagation discrepancy.
$|\Delta r_{\mathrm{prop}}|$ (deg) $|\Delta p_{\mathrm{prop}}|$ (deg) $\Delta\phi$ (deg)
Trial 1 0.18 $\pm$ 0.24 0.13 $\pm$ 0.14 0.22 $\pm$ 0.21
Trial 2 0.20 $\pm$ 0.26 0.24 $\pm$ 0.38 0.31 $\pm$ 0.35
Trial 3 0.11 $\pm$ 0.13 0.14 $\pm$ 0.14 0.18 $\pm$ 0.14
Mean 0.16 $\pm$ 0.21 0.17 $\pm$ 0.22 0.24 $\pm$ 0.23
Note: Per-step increments are $\Delta R_{i}=R_{i-1}^{\top}R_{i}$ and discrepancy is $\Delta R^{\mathrm{diff}}_{i}=(\Delta R^{O}_{i})^{\top}\Delta R^{V}_{i}$. Yaw is constrained (shaft spin locked) in the estimator, so we report roll/pitch and the geodesic discrepancy angle $\Delta\phi$.
Discussion.
Across three trials, the proposed vision-based method achieved millimeter-scale agreement with optical tracking in the anatomy frame (Table I) while maintaining low inter-frame rotation propagation discrepancy (Table II). The largest translation deviations occur primarily during fast motion and when the tool undergoes large out-of-plane tilt, which is consistent with increased depth sensitivity and partial visibility.
IV-B Workflow Efficiency
We evaluated workflow efficiency by measuring the user time required for (1) interactive tool segmentation and mask selection and (2) anatomy registration. Each experiment was repeated across multiple trials performed by the same user under consistent conditions. Table III summarizes the recorded times.
On average, tool segmentation and selection required 0 min $27\pm 2$ sec, anatomy registration required $1$ min $15\pm 15$ sec, and the complete workflow was completed in $1$ min $42\pm 12$ sec.
@@ IV-B Workflow Efficiency
IV-C Case Study on Live Video Stream
To assess usability under realistic operating conditions, we performed a mock surgical study on a live endoscopic video stream in which the user executed the complete pipeline end-to-end: tool segmentation during continuous motion, temporal mask propagation, anatomy segmentation and refinement, anatomy registration, and depth-aware anatomical overlay. Figure 11 shows representative snapshots of the interactive workflow and the resulting registered overlay.
Qualitatively, the system sustained stable tool tracking under moderate occlusions. The on-demand agent interface also allowed the user to transition quickly between segmentation, refinement, and registration without interrupting sterile workflow. Overall, these observations suggest the feasibility of speech-guided, vision-based navigation support for skull-base procedures with minimal setup overhead.
@@ V Conclusion
In this work, we presented an embodied, speech-guided agent framework for video-guided skull base surgery that integrates natural language interaction with real-time visual perception and image-guided navigation. The proposed architecture enables a modular workflow by decoupling reasoning from perception, in which a language-driven planner interprets surgeon intent while specialist vision models execute segmentation, tracking, registration, and pose estimation directly on live endoscopic video streams.
Building upon the prior SCOPE framework, which focused on speech-guided segmentation and tracking, the present work extends the system toward a unified task manager capable of orchestrating multiple perception and navigation functions. The framework enables interactive tool and anatomy segmentation, tool pose estimation, registration, and depth-aware overlays, thereby allowing surgeons to trigger image-guided workflows through natural language queries. Quantitative evaluation demonstrates that the proposed vision-based approach achieves competitive spatial accuracy compared with a commercial optical tracking system while reducing hardware dependence and simplifying workflow integration.
Despite promising results, this study primarily demonstrates technical feasibility in controlled experimental settings, including mock surgical scenarios. The next next step includes evaluating the utility of the system with surgeons. This will include assessing both technical accuracy and task-level outcomes, including time-to-completion, error rates, and interaction efficiency. Additionally, we will incoporate human-centered metrics such as cognitive workload, usability, and perceived workflow compatibility, which will ultimately determine whether the system can effectively function as an intraoperative surgical assistant. Future work will also explore extension to other imaging modalities and surgical procedures. Ultimately, by combining language-guided interaction with real-time surgical perception, the proposed framework represents a step toward scalable, hardware-light intelligent assistants that support surgeons in complex image-guided interventions.

### 16 Apr 2026 v2

@@ IV-B Pose Tracking Accuracy
TABLE II: Tool-tip translation, inter-frame rotation propagation errors, and runtime for depth-based and hybrid pose estimation methods.
Translation error in camera frame Rotation propagation discrepancy in camera frame
Method Trial Speed (FPS) $|\Delta x|$ ($m$) $|\Delta y|$ ($m$) $|\Delta z|$ ($m$) $\|\Delta\mathbf{p}\|_{2}$ ($m$) $|\Delta y_{\mathrm{prop}}|$ (deg) $|\Delta p_{\mathrm{prop}}|$ (deg) $\Delta\phi$ (deg)
Depth Anything Trial 1 8.13 3.59 $\pm$ 0.76 0.80 $\pm$ 0.70 24.03 $\pm$ 5.99 24.35 $\pm$ 5.94 0.27 $\pm$ 0.37 10.88 $\pm$ 9.78 10.91 $\pm$ 9.76
Trial 2 8.21 0.42 $\pm$ 0.44 0.60 $\pm$ 0.42 11.77 $\pm$ 7.17 11.85 $\pm$ 7.11 0.33 $\pm$ 0.41 10.34 $\pm$ 9.40 10.37 $\pm$ 9.38
Trial 3 8.03 2.79 $\pm$ 0.87 1.20 $\pm$ 0.51 12.82 $\pm$ 7.20 13.36 $\pm$ 6.91 0.60 $\pm$ 4.85 8.58 $\pm$ 9.13 8.67 $\pm$ 9.37
Mean 8.12 2.267 $\pm$ 0.66 0.87 $\pm$ 0.54 16.21 $\pm$ 6.78 16.52 $\pm$ 6.65 0.40 $\pm$ 1.88 9.93 $\pm$ 9.44 9.98 $\pm$ 9.50
Video Depth Anything Trial 1 6.20 3.73 $\pm$ 0.73 0.78 $\pm$ 0.65 26.90 $\pm$ 5.74 27.19 $\pm$ 5.72 0.23 $\pm$ 0.31 10.78 $\pm$ 10.38 10.81 $\pm$ 10.36
Trial 2 6.12 0.47 $\pm$ 0.44 0.67 $\pm$ 0.45 15.54 $\pm$ 9.45 15.61 $\pm$ 9.39 0.26 $\pm$ 0.36 12.98 $\pm$ 14.37 13.02 $\pm$ 14.34
Trial 3 6.31 3.39 $\pm$ 0.85 1.18 $\pm$ 0.54 24.47 $\pm$ 7.73 22.5 $\pm$ 7.96 0.31 $\pm$ 0.40 0.21 $\pm$ 11.46 11.83 $\pm$ 11.80
Mean 6.21 2.53 $\pm$ 0.67 0.87 $\pm$ 0.55 22.30 $\pm$ 6.78 21.77 $\pm$ 7.69 0.27 $\pm$ 0.36 7.99 $\pm$ 12.07 11.89 $\pm$ 12.17
Hybrid Approach Trial 1 9.98 2.15 $\pm$ 0.73 0.87 $\pm$ 0.58 0.74 $\pm$ 0.75 2.55 $\pm$ 0.92 0.13 $\pm$ 0.14 0.17 $\pm$ 0.23 0.29 $\pm$ 0.27
Trial 2 10.11 0.39 $\pm$ 0.40 0.45 $\pm$ 0.36 1.50 $\pm$ 1.84 1.76 $\pm$ 1.78 0.20 $\pm$ 0.27 0.26 $\pm$ 0.39 0.31 $\pm$ 0.45
Trial 3 10.05 2.14 $\pm$ 0.59 1.14 $\pm$ 0.46 0.78 $\pm$ 0.59 2.66 $\pm$ 0.60 0.20 $\pm$ 0.33 0.21 $\pm$ 0.29 0.40 $\pm$ 0.44
Mean 10.05 1.56 $\pm$ 0.57 0.82 $\pm$ 0.47 1.01 $\pm$ 1.06 2.32 $\pm$ 1.10 0.18 $\pm$ 0.25 0.21 $\pm$ 0.30 0.37 $\pm$ 0.39
Note: Translation differences are between the optical tracker and the vision-based estimate, with all quantities expressed in the camera frame. Rotation discrepancy is computed from inter-frame camera-frame increments, where $\Delta R_{i}=R_{i-1}^{\top}R_{i}$ and $\Delta R_{i}^{\mathrm{diff}}=(\Delta R_{i}^{O})^{\top}\Delta R_{i}^{V}$. Because roll is constrained in the estimator, we report only yaw, pitch, and the corresponding geodesic discrepancy angle $\Delta\phi$, which is also computed from yaw and pitch only. Values are mean $\pm$ standard deviation over time-matched frames for each trial. Speed is reported as frames per second (FPS) on a single NVIDIA RTX 4090 GPU, measured for pose tracking inference only without visualization, and summarized in the mean row.
Inter-frame rotation propagation discrepancy. Let $\mathbf{R}^{V}_{C,i}\in SO(3)$ and $\mathbf{R}^{O}_{C,i}\in SO(3)$ denote the vision-based and optical-tracker rotations expressed in the camera frame. We compute the inter-frame increments
$\Delta\mathbf{R}^{(\cdot)}_{i}=\left(\mathbf{R}^{(\cdot)}_{C,i-1}\right)^{\!\top}\mathbf{R}^{(\cdot)}_{C,i}$, (4)
and define the propagation discrepancy as
@@
Figure 10: Trial 1 tool-tip trajectory in the camera frame (Hybrid Approach). The $x$, $y$, and $z$ components are plotted over time (time-matched sample index) for the vision-based method (Blue) and the optical tracker (Orange).
Figure 11: Trial 2 tool-tip trajectory in the camera frame (Hybrid Approach).
Figure 12: Trial 3 tool-tip trajectory in the camera frame (Hybrid Approach).
The results show a clear advantage for the proposed hybrid formulation. In particular, the depth-only variants exhibited large errors along the camera depth axis and substantially higher rotation propagation discrepancy. In contrast, the hybrid method achieved a mean Euclidean tip-position error of $2.32\pm 1.10$ mm, with sub-degree inter-frame rotation discrepancy, while also operating at approximately 10 FPS. These results indicate that combining registration-grounded depth cues with mask-based geometric constraints can result in stable pose tracking in this setting. Figures 1012 visualize the tool-tip trajectories in the camera frame for the hybrid method and optical tracker across all three trials.
IV-C Workflow Efficiency
In addition to spatial accuracy, we evaluated the practical efficiency of the interactive workflow by measuring the time required for: (1) tool segmentation and mask selection, and (2) anatomy registration. Each experiment was repeated across three trials performed by the same user under consistent conditions. Table III summarizes the recorded times.
Overall, interactive tool segmentation and selection required $26\pm 1$ sec, anatomy registration required $1$ min $22\pm 5$ sec, and the full workflow was completed in $1$ min $48\pm 5$ sec. These results indicate that the proposed interaction design supports rapid setup and may be practical for intraoperative use without introducing major workflow overhead.
@@ IV-C Workflow Efficiency
IV-D Discussion
Overall, the results demonstrate that the proposed system performs well across two complementary dimensions: accuracy, and workflow efficiency. The full workflow was completed in under two minutes across all three trails, indicating that the interaction design is efficient and practical for intraoperative use without introducing significant overload.
The registration results show that the virtual cursor achieves sub-millimeter accuracy and remains comparable to manual clicking, despite relying entirely on vision-based tool-tip localization. This highlights the reliability of the underlying 2D tip tracking formulation and supports the feasibility of hands-free interaction as a core component of the system. Since the same tip estimation is used across multiple stages—including registration, anatomy segmentation, and pose tracking—these results validate it as a stable and consistent building block of our system.
The pose tracking results show the depth-derived formulation provides a useful geometric estimate, but it does not explicitly enforce temporal consistency across frames. Consequently, the recovered tool position along the camera depth axis exhibits noticeable frame-to-frame fluctuations, even when the underlying motion is relatively smooth. We also evaluated Video Depth Anything, which would be expected to improve temporal coherence for sequential data. However, in our experiments, this was not the case: the resulting depth trajectories showed even larger excursions and more pronounced outliers. These findings indicate that monocular depth alone is insufficient for stable surgical tool pose tracking.
Figure 13: Tool-tip depth trajectory in the camera frame using the depth-derived formulation. The estimated tool-tip position along the camera $z$ axis shows substantial frame-to-frame fluctuation relative to the optical tracker reference. While the coarse trend is preserved, the trajectory contains frequent abrupt variations, reflecting limited temporal consistency when pose is inferred directly from monocular depth.
Figure 14: Tool-tip depth trajectory in the camera frame using Video Depth Anything. Although designed for video input, Video Depth Anything does not improve stability in this setting. The estimated depth trajectory still contains substantial fluctuations and, in some cases, even larger deviations than the frame-wise depth-derived approach.
These limitations motivated the proposed hybrid formulation, which combines registration-grounded depth cues with image-plane geometric constraints to improve temporal stability. Specifically, the relative monocular depth prediction is converted to metric depth using the depth prior obtained from the registered anatomy. This provides a scene-consistent 3D reference in the camera frame for recovering coarse tool structure, while the observed 2D mask geometry constrains the tool orientation and projected extent. By combining anatomy-anchored metric depth with mask-based geometric constraints, the hybrid approach suppresses the large outliers seen in depth-only methods and produces smoother, more reliable pose estimates over time.
The advantage of the vision-based pipeline lies not only in reducing reliance on additional external tracking equipment, but also in providing greater robustness to failure modes that commonly affect such systems. In particular, optical trackers may suffer from intermittent dropouts due to line-of-sight constraints. In our experiments, these failures occasionally appeared as abrupt spikes in the measured rotation propagation, reflecting limitations of the external reference system rather than instability in the proposed method.
Figure 15: Example of per-step roll propagation in the camera frame. Absolute inter-frame roll changes are shown for the proposed vision-based tracker and the optical tracker. The vision-based estimate remains smooth over time, while the optical tracker exhibits an isolated spike, likely caused by transient line-of-sight loss or marker occlusion. This example highlights the robustness of the proposed method to common failure modes of external tracking systems.
Taken together, these results suggest that the proposed system formulation not only achieves competitive accuracy relative to optical tracking, but also provides improved robustness and workflow integration. This positions vision-based, speech-guided interaction as a viable alternative for real-time surgical navigation and guidance.
@@ V Conclusion
In this work, we presented an embodied, speech-guided agent framework for video-guided skull base surgery that integrates natural language interaction with real-time visual perception and image-guided navigation.
By decoupling reasoning from perception, the proposed architecture supports a modular workflow in which a language-driven planner interprets surgeon intent while specialist vision models perform segmentation, tracking, registration, and pose estimation directly on live endoscopic video streams. This design allows the system to integrate multiple vision capabilities within a single interactive framework while avoiding the complexity of tightly coupled end-to-end multimodal models.
The framework enables interactive tool segmentation, anatomy segmentation, tool pose estimation, anatomy registration, and depth-aware anatomical overlays, thereby allowing surgeons to trigger image-guided workflows through natural language commands. Across quantitative experiments, the proposed vision-based pipeline achieved competitive spatial accuracy relative to a commercial optical tracking system while reducing reliance on additional tracking hardware and supporting a streamlined, hands-free workflow.
Despite promising results, this study primarily demonstrates technical feasibility in controlled experimental settings, including mock surgical scenarios.
The next next step includes evaluating the utility of the system with surgeons. This will include assessing both technical accuracy and task-level outcomes, including time-to-completion, error rates, and interaction efficiency.
Additionally, we will incoporate human-centered metrics such as cognitive workload, usability, and perceived workflow compatibility, which will ultimately determine whether the system can effectively function as an intraoperative surgical assistant.
Future work will also explore extension to other imaging modalities and surgical procedures. Ultimately, by combining language-guided interaction with real-time surgical perception and navigation, the proposed framework represents a step toward scalable, hardware-light intelligent assistants that support surgeons in complex image-guided interventions.